OpenStudy (anonymous):

A bag contains a total of 20 batteries, of which six are defective. Selecting two at random, without replacement, determine the probability that none of the batteries you select are good. 3/5 3/10 3/38 9/100

OpenStudy (anonymous):

\[\frac{14}{20}\times \frac{13}{19}\]

OpenStudy (smurfy14):

6/20 times 5/19

OpenStudy (anonymous):

i get \[\frac{91}{190}\]

OpenStudy (smurfy14):

which equals 30/380..now reduce

OpenStudy (anonymous):

thats what i got satellite

OpenStudy (anonymous):

oh i am wrong. i thought it said "none are defective" it says "none are good" smurfy is right

OpenStudy (smurfy14):

it simplifies to 3/38 so it is the third one

OpenStudy (anonymous):

i wonder why it does not say "both are defective" to make life easier

OpenStudy (anonymous):

lol cuz this class sucks... lmao